OpenStudy (aaronandyson):

A particle moving with a constant acceleration describes in the last second of it motion 9/25th of the whole dist. if it starts from rest,how long is the particle is the particle in motion and through what dist does it movie if it describes 6cm in the first sec?

OpenStudy (phi):

is this calculus ?

OpenStudy (aaronandyson):

Kinematics.

OpenStudy (aaronandyson):

And yes,I did learn some Calculus.

OpenStudy (phi):

do you have a formula for distance as a function of acceleration ?

OpenStudy (aaronandyson):

S = ut + 0.5at^2

OpenStudy (phi):

ok, and they say it starts from rest that tells us what u is. any idea ?

OpenStudy (aaronandyson):

If you the body is at rest ,then u = 0 ms^-1.

OpenStudy (aaronandyson):

Therefore,S = 0.5at^2

OpenStudy (phi):

sounds good. it describes 6cm in the first sec if we use that info, what do we get for a ?

OpenStudy (aaronandyson):

Idk. Thats where I'm stuck. I have no idea to move on any further.

OpenStudy (phi):

you know this S = 0.5at^2 you know it describes 6cm in the first sec which you have to interpret: it moved how far? how long did it take ? that info corresponds to your equation: s and t.

OpenStudy (aaronandyson):

I'm so confused as idk what to do.. e.e...

OpenStudy (aaronandyson):

do i need to take the distance travaled as 9/25 in t-1 seconds?

OpenStudy (phi):

"describe" is not the clearest word (the 5th or 6th definition is "trace the outline" I would use the word "traveled"

OpenStudy (phi):

right now, we are concentrating on these two things: s = 0.5 a t^2 and it traveled 6 cm in the first second.

OpenStudy (aaronandyson):

so then 0.06 = 0.5*a*1 a = 0.012ms^-2?

OpenStudy (phi):

yes though maybe it's easier to work in cm rather than meters to avoid the decimals we can switch to meters at the end if we have to.

OpenStudy (phi):

now we know what a is s= 0.5 * a * t^2 s= 6 t^2 (in cm , not meters)

OpenStudy (aaronandyson):

it should be 6*10^-2 m?

OpenStudy (phi):

yes, it can be written in meters. In that case, s= 0.06 t^2 (in meters)

OpenStudy (aaronandyson):

why didnt you include a in the equation?

OpenStudy (phi):

we started with s = 0.5 a t^2 we put in 0.06 m for s, and t=1: 0.06 m = 0.5 *a *1 sec^2 mulitply both sides by 2 0.12 m/sec^2 = a

OpenStudy (phi):

now that we know a, we can write s= 0.5 * a * t^2 and put in a; s= 0.5*0.12*t^2 and simplify that to s = 0.06 t^2 that is the formula for how far the particle moves as a check: after 1 second we get s= 0.06 * 1 = 0.06 m= 6 cm which is what it should be. ok ?

OpenStudy (aaronandyson):

okay.

OpenStudy (aaronandyson):

What next?

OpenStudy (phi):

we want to find how long is the particle is the particle in motion which means find "t" (how long it moved) we have to use the last info they gave us. describes in the last second of it motion 9/25th of the whole dist.

OpenStudy (aaronandyson):

so when t-1 then s = 0.5a(t-1)^2?

OpenStudy (phi):

I'm wondering what is the nicest way to to think about it. we know a few things if we let the total amount of time be T then s(T) - s(T-1) = (9/25) s(T) s(T) is short for "how far it moved in T seconds"

OpenStudy (aaronandyson):

s(t-1) = 9/25 of S

OpenStudy (phi):

no. it moves in the last second of it motion 9/25th of the whole distance the "whole distance" is s(T) if it moved 9/25 of the distance in the last second, that means it moved 16/25 of the distance in the time from 0 up to T-1 in other words, the distance it moves in time T-1 is 16/25 of the entire distance s(T-1) = 16/25 s(T)

OpenStudy (phi):

s(T-1) = 16/25 s(T) or 25 s(T-1) = 16 s(T)

OpenStudy (aaronandyson):

howd you get 15?typo?

OpenStudy (phi):

somehow I changed 25 into 15. gack!

OpenStudy (phi):

can you solve the "quadratic equation" that we get ?

OpenStudy (aaronandyson):

25 s(T-1) = 16 s(T) is not an quadratic eqn....

OpenStudy (aaronandyson):

25(t^2 + 2t+1) = 16

OpenStudy (phi):

remember s(T) is short for 0.06*T^2 and s(T-1) is short for 0.06*(T-1)^2

OpenStudy (aaronandyson):

25*0.06*(t-1)^2 = 16*0.06*t^2 so 0.06 get cancelled?

OpenStudy (phi):

yes the 0.06 cancels we have to expand (t-1)^2 collect terms

OpenStudy (aaronandyson):

t^2 -2t + 1

OpenStudy (phi):

ok, put that into the equation to replace (t-1)^2

OpenStudy (aaronandyson):

25t^2 -50t + 25

OpenStudy (phi):

and don't forget the other side

OpenStudy (aaronandyson):

9t^2 - 50t + 25 = 0

OpenStudy (phi):

I used the quadratic formula, but it does factor.

OpenStudy (aaronandyson):

i got 5/9 and 5

OpenStudy (phi):

that is the time in seconds. we know it goes at least 1 second, so the first solution 5/9 is not relevant

OpenStudy (aaronandyson):

t = 5s

OpenStudy (phi):

and now you find the total distance s = 0.06 t^2

OpenStudy (aaronandyson):

1.5 m

OpenStudy (aaronandyson):

oops...1.5 cm

OpenStudy (phi):

no , it's meters remember we are using 0.06 m/sec^2

OpenStudy (aaronandyson):

so is the final ans 1.5 m?

OpenStudy (phi):

and also it describes 6cm in the first sec? so you know the answer can't be 1.5 cm (which is smaller than 6 cm) yes, the answer is it takes 5 seconds, and travels a total distance of 1.5 meters

OpenStudy (aaronandyson):

ty